Graham Petroleum produces oil. On May 1, it had no work-in-process inventory. It started production of 215 million barrels of oil in May and shipped 185 million barrels in the pipeline. The costs of the resources used by Graham in May consist of the following:

Materials $ 2,532 million
Conversion costs (labor and overhead) $ 3,320 million


Required:

The production supervisor estimates that the ending work in process is 80 percent complete on May 31. Compute the cost of oil shipped in the pipeline and the amount in work-in-process ending inventory as of May 31. (Do not round intermediate calculations. Enter your answers in millions. For example, enter "1" instead of "1,000,000".)

Cost of Oil shipped in the pipeline? (million)

Work in Process ending inventory? (million)
Answer #1

Completed barrels = 185 + (215 - 185) X 80% = 209

Cost per barrel = ($2,532 + $ 3,320) / 209 = $28

Cost of Oil shipped in the pipeline = 185 X $28 = $5,180

Work in Process ending inventory = (215 - 185) X 80% X $28 = $672